OpenStudy (anonymous):

How do you find all the solutions for x in the equation x+ <22-3x> =6 (Everything in <> is square rooted)

OpenStudy (anonymous):

\[\implies \sqrt{22-3x} = 6-x\]\[\implies 22-3x = (6-x)^2; \qquad x \le \frac{22}{3}\]\[\implies 22 - 3x = 36 - 12x + x^2; \qquad x \le \frac{22}{3}\]\[\implies x^2 - 9x + 14 = 0; \qquad x \le \frac{22}{3}\]\[\implies (x-7)(x-2) = 0; \qquad x \le \frac{22}{3}\]\[\implies x \in \{2,7\}\]

myininaya (myininaya):

we all want 6-x>0 6>x or x<6

OpenStudy (anonymous):

You can disregard the x=7 solution though, because if you substitute 7 back in the original problem: 7 + √(22 - 3(7)) = 6 7 + √(22 - 21) = 6 7 + √(1) = 6 7 + 1 = 6

myininaya (myininaya):

also*

myininaya (myininaya):

we want 6-x>0 since sqrt(something) has domain something>0

myininaya (myininaya):

something >=0

myininaya (myininaya):

6-x>=0 x<=6

myininaya (myininaya):

so just 2 is what i'm saying

myininaya (myininaya):

since that is only element that was less than or equal to 6

OpenStudy (anonymous):

Oh, good point. Forgot to keep track of that restriction.